# Reconstructing Quantitative Cerebral Perfusion Images Directly From Measured Sinogram Data Acquired Using C-arm Cone-Beam CT

### v1

@@ IV-C Quantitative Assessment of Perfusion Parametric Images
The comparison of reconstruction error quantified by overall nrMSE and PSNR using baseline methods and TRAINER was shown in Tab. I and Tab. II. Compared to all baseline methods, TRAINER achieves the highest overall quantitative accuracy at all exposure levels. Additionally, we conducted a detailed statistical analysis on ten selected regions of interest (ROIs). The results demonstrated that in MTT for healthy tissue, penumbra, and ischemic core, the PSNR metrics were significantly improved under four different exposure levels compared to the best performance of the baseline methods. Specifically, the PSNR for healthy tissue increased by an average of 11.7%, with a particularly notable increase in the penumbra region, reaching an average of 70.6%, and an average increase of 36.6% in the core region. Similarly, a comparable trend was observed in the analysis of CBF. For CBF images in healthy tissue, penumbra, and core, the PSNR achieved an average increase of 20.22%, 21.16%, and 33.19% under the four exposure levels, respectively, compared to the best performance of the baseline methods. These findings underscore the effectiveness of TRAINER in improving quantitative accuracy and image quality.
TABLE I: The comparison of reconstruction error quantified by overall nrMSE using baseline methods and TRAINER. Data were simulated at gantry rotation speed of $dt=8$ sec and four exposure levels.
@@
CBF Fbp+Tikh 27.84% 27.85% 27.84% 29.17%
Fbp+TTV 26.79% 26.95% 27.62% 32.93%
TRT+TTV 22.89% 26.23% 28.67% 41.06%
eSMART+TTV 19.62% 28.80% 27.53% 39.38%
TRAINER 9.07% 13.50% 13.98% 22.34%
MTT Fbp+Tikh 46.99% 52.78% 52.6% 57.67%
Fbp+TTV 51.63% 37.53% 38.54% 44.82%
TRT+TTV 20.37% 35.80% 36.02% 42.19%
eSMART+TTV 27.21% 25.6% 29.74% 43.18%
TRAINER 18.28% 20.86% 20.92% 23.59%
@@
MTT Fbp+Tikh 27.66 27.67 27.69 27.79
Fbp+TTV 26.85 26.89 26.96 27.19
TRT+TTV 33.58 32.93 32.69 31.76
eSMART+TTV 30.96 31.44 31.37 30.94
TRAINER 45.16 39.31 41.32 37.89
The comparison of reconstruction error quantified by PSNR for each tissue types (healthy tissue, penumbra, ischemic core) using baseline methods and TRAINER was shown in Fig. 6 and Fig. 7. Data were simulated at gantry rotation speed of $dt=8$ sec and four exposure levels. Compared to the baseline methods, TRAINER improves the quantitative accuracy of perfusion parametric images for all types of tissues. PSNR values were calculated by carefully selecting ten regions of interest.
Figure 6: The comparison of reconstruction error of CBF quantified by tissue-specific PSNR using baseline methods and TRAINER. Data were simulated at gantry rotation speed of $dt=8$ sec and four exposure levels.
Figure 7: The comparison of reconstruction error of MTT quantified by tissue-specific PSNR using baseline methods and TRAINER. Data were simulated at gantry rotation speed of $dt=8$ sec and four exposure levels.
@@ V-A Potential Limitations and Future Work
Second, the subject-specific generative model in Eq. (11) is conditioned by its input, $\mathbf{Z}$. Currently, an empirical selection of $\mathbf{Z}$ has been employed to generate results shown in the paper. It is very likely that the quantitative accuracy and image quality of the reconstructed perfusion parametric images depend on the selection of $\mathbf{Z}$. In our future studies, we will investigate the performance dependence on different selections such as time-resolved FBP images. In addition, the condition $\mathbf{Z}$ may not necessarily be treated as a constant, instead, it can be treated as a argument to be optimized under the ADMM iterative framework. In our future studies, we will also explore the technical feasibility of optimizing the condition and investigate the potential benefits of this novel strategy.

### v2

@@ IV-C Quantitative Assessment of Perfusion Parametric Images
The comparison of reconstruction error quantified by overall rRMSE and PSNR using baseline methods and TRAINER is shown in Tab. I and Tab. II. Compared to all baseline methods, TRAINER achieves the highest overall quantitative accuracy at all exposure levels. Additionally, we conducted a detailed statistical analysis on ten selected regions of interest (ROIs). The results demonstrated that in MTT for healthy tissue, penumbra, and ischemic core, the rRMSE metrics were significantly improved under four different exposure levels compared to the best performance of the baseline methods. Specifically, the rRMSE for healthy tissue decreased by an average of $60.56\%$, with a particularly notable decrease of $87.40\%$ in the penumbra region and a $72.45\%$ decrease in the ischemic core region. For CBF images in healthy tissue, penumbra, and ischemic core, the rRMSE decreased by averages of $42.22\%$, $15.26\%$, and $8.23\%$, respectively, under the four exposure levels, compared to the best performance of the baseline methods. These findings underscore the effectiveness of TRAINER in improving quantitative accuracy and image quality.
TABLE I: The comparison of reconstruction error quantified by overall rRMSE using baseline methods and TRAINER. Data were simulated at gantry rotation speed of $dt=8$ sec and four exposure levels.
@@
CBF Fbp+Tikh 39.60% 39.61% 39.60% 39.68%
Fbp+TTV 34.59% 34.28% 34.32% 34.54%
TRT+TTV 12.97% 13.64% 14.62% 18.14%
eSMART+TTV 13.49% 15.16% 15.74% 20.14%
TRAINER 4.94% 6.11% 7.24% 10.46%
MTT Fbp+Tikh 32.18% 32.15% 32.08% 31.69%
Fbp+TTV 35.10% 34.96% 34.69% 33.84%
TRT+TTV 18.64% 19.30% 19.34% 20.08%
eSMART+TTV 23.57% 22.50% 22.56% 23.24%
TRAINER 3.60% 7.08% 5.63% 8.31%
@@
MTT Fbp+Tikh 26.13 26.13 26.15 26.26
Fbp+TTV 25.37 25.41 25.47 25.69
TRT+TTV 30.87 30.57 30.55 30.22
eSMART+TTV 28.83 29.24 29.21 28.95
TRAINER 45.16 39.27 41.26 37.89
The comparison of reconstruction error quantified by rRMSE for each tissue type (healthy tissue, penumbra, ischemic core) using baseline methods and TRAINER was shown in Fig. 6 and Fig. 7. Data were simulated at gantry rotation speed of $dt=8$ sec and four exposure levels. Compared to the baseline methods, TRAINER improves the quantitative accuracy of perfusion parametric images for all types of tissues. rRMSE values were calculated by carefully selecting ten regions of interest.
Figure 6: The comparison of reconstruction error of CBF quantified by tissue-specific rRMSE using baseline methods and TRAINER. Data were simulated at gantry rotation speed of $dt=8$ sec and four exposure levels.
Figure 7: The comparison of reconstruction error of MTT quantified by tissue-specific rRMSE using baseline methods and TRAINER. Data were simulated at gantry rotation speed of $dt=8$ sec and four exposure levels.
@@ V-A Potential Limitations and Future Work
Second, the subject-specific generative model in Eq. (13) is conditioned by its input, $\mathbf{Z}$. Currently, an empirical selection of $\mathbf{Z}$ has been employed to generate results shown in the paper. It is very likely that the quantitative accuracy and image quality of the reconstructed perfusion parametric images depend on the selection of $\mathbf{Z}$. In our future studies, we will investigate the performance dependence on different selections such as time-resolved FBP images. In addition, the condition $\mathbf{Z}$ may not necessarily be treated as a constant, instead, it can be treated as an argument to be optimized under the ADMM iterative framework. In our future studies, we will also explore the technical feasibility of optimizing the condition and investigate the potential benefits of this novel strategy.